### v1

@@ 6.1 General Approach
In the context of mean-field control or games, an alternative connection to an optimization problem is proposed for several models, including [14, 60, 61] for the Kuramoto synchronization and [54, 55] for the Cucker-Smale flocking. In general, we propose the following MFC with a separable structure,
@@
This is again an equation of type (1.1), except the functional $F$ is replaced by the value function $v$. Now, a standard control argument shows that linear convexity of $F$ implies linear convexity of the value function $v$, suggesting a similar structure of (1.1) and (6.2). Further, note that if $F$ is sufficiently smooth, one can always define a suitable control problem such that the optimally controlled state dynamics exactly match (1.1). Indeed, this can be achieved by replacing $F$ by,
$\widetilde{F}(\mu):=\beta F(\mu)+\frac{1}{2}\mu\left(|\nabla_{x}\delta_{\mu}F(\mu)(\cdot)|^{2}\right)-\frac{1}{2}\mu\left(\delta_{\mu}F(\mu)(\cdot)\right).$
Then, a direct calculation shows that the value function with running cost $\widetilde{F}$ is exactly $F$, and the optimal state process solves the mean-field Langevin equation (1.1).
@@
and state dynamics $X^{\alpha}_{t}=X_{0}+\int_{0}^{t}\alpha_{s}\,\mathrm{d}s+\sigma W_{t}$. We continue to illustrating the above discussion with two prominent examples of mean-field games.
@@ 6.2 Kuramoto Synchronization
An intruging property of this model is a phase transition across a critical value of $\kappa$, from incoherence to spontaneous synchronization. The related mean-field model studied in [28] that shares this feature is a McKean-Vlasov equation obtained by setting $\omega^{k}$ to zero and adding a Brownian noise with strength $\sigma>0$. The resulting dynamics read,
@@
with the one-dimensional torus $\mathbb{T}$ as its state space. This flow is in the form of a mean-field Langevin SDE (1.1), for the potential
@@
subject to $X^{\alpha}_{t}=\int_{0}^{t}\alpha_{s}\,\mathrm{d}s+\sigma W_{t}$. In view of Theorem 3.6, any minimizer $\alpha^{*}$ of this MFC gives rise to a mean-field Nash equilibrium $\mu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{t})$, in which players minimize,
@@
The above MFG approach to synchronization was first proposed by Yin, Mehta, Meyn, & Shanbhag in [61, 60], and later used by Carmona & Graves [12] to study jet-lag recovery by modeling the alignment with the circadian rhythm. These studies and the recent works of Carmona, Cormier, & Soner [14], Cesaroni & Cirant [18], and Höfer & Soner [33] establish phase transitions in the Kuramoto MFGs analogous to the one exhibited by the original and mean-field Kuramoto dynamical system, providing evidence for a connection between these seemingly very different models.
@@ 6.3 Cucker-Smale Flocking
Flocking is interpreted as $\max_{k,j=1,..,N}|x^{k}_{t}-x^{j}_{t}|$ staying bounded for all times and [21] proves that flocking occurs if $\rho<1/2$. Motivated by this model, Nourian, Caines & Malhamé [54, 55] propose an ergodic mean-field game in which the representative player’s running cost is given by,
@@
Consistent this observation, [11, Chapter 4.7.3] suggests a slightly different MFG than that of [54]. It is a potential game corresponding to the MFC of (6.1) with running cost $F$ and the $2n$-dimensional state dynamics
@@ 6.4 Linear Quadratic Problem
The special case $\rho=0$ of the Cucker-Smale model is the well-known linear quadratic problem, which we now discuss. In this case, the mean-field limit potential is given by,
$F(\mu)=\frac{1}{2}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}|x-y|^{2}\,\mu(\mathrm{d}x)\,\mu(\mathrm{d}y)=\frac{1}{2}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}(|x|^{2}+|y|^{2}-2x.y)\,\mu(\mathrm{d}x)\mu(\mathrm{d}y)=\mathrm{Var}(\mu),$
@@
This implies that the corresponding mean-field Langevin equation is given by,
@@
The stationary measures of this equation are Gaussian distributions with arbitrary mean and variance $\sigma^{2}/(4a)$. The associated potential mean-field game minimizes
$J_{g,\beta}(X_{0},\alpha,\mu)=\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}|\alpha_{t}|^{2}+\kappa\int_{\mathbb{R}}|X^{\alpha}_{t}-y|^{2}\,\mu_{t}(\mathrm{d}y)\right)\,\mathrm{d}t.$
for a given continuous measure flow $t\in[0,\infty)\mapsto\mu_{t}\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$. This mean-field game can be solved explicitly, and a direct computation shows that any Gaussian distribution with variance $\sigma^{2}/(4a)$ provides a stationary mean-field game equilibrium.
@@ 7 Invariance Principle
In this section, we prove the invariance principle stated in Definition 3.2 under under the following regularity and growth conditions. Let $\mathfrak{m}_{2}$ be the 2-Wasserstein metric on ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ and set ${\mathscr{X}}:={\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$. Recall the sup-norm $\|\cdot\|_{T}$ on ${\mathscr{D}}_{n}$, the notation ${\mathfrak{s}}(\nu)$ for the second moment (3.5), and the value function $v(X_{0},\gamma)$ of (3.1).
@@
In addition Assumption 2.3, the followings hold:
  • (Local Uniform Boundedness). For any $\ell>0$, there exists a constant $c_{\ell}$ such that for any $(t,x,\zeta)\in[0,T]\times{\mathscr{D}}_{n}\times E$ and $|a|,|a^{\prime}|\leq\ell$,
    $|(b,\sigma,\sigma^{0})(t,x,a)-(b,\sigma,\sigma^{0})(t,x,a^{\prime})|$ $\leq c_{\ell},$
    $|(\lambda,\lambda^{0})(t,x,a,\zeta)-(\lambda,\lambda^{0})(t,x,a^{\prime},\zeta)|$ $\leq c_{\ell}L(\zeta).$
@@
  • (Uniform Continuity). For any $K>0$, there is a continuous map $m_{K}:[0,\infty)\mapsto[0,4K]$ with $m_{K}(0)=0$ such that for all $(t,x,\hat{x},\nu,\hat{\nu},a,\hat{a})\in[0,T]\times{\mathscr{X}}^{2}$,
@@
If the coefficients of an MFC problem satisfy Assumptions 2.3 and 7.1, then it is invariant under probability bases.
@@
In the diffusion setting, the invariance is proved in [20, Theorem 3.6] for ‘rich enough’ probabilistic structures [20, Section 2.1.3] and in [26, Proposition 2.4]. The counter example provided in [20, Section 3] shows the necessity of continuity assumptions. A careful analysis of the proof of our results shows that local uniform boundedness assumption can be replaced by ${\mathscr{L}}^{p}$ integrability of $X_{0}$ for some $p>2$, a condition used in several places, for example [44], to overcome similar difficulties.
This section is devoted to the proof of the above Theorem. Let ${\mathscr{C}}_{m}$ be the set of $\mathbb{R}^{m}$ valued continuous functions on $[0,T]$, and recall the measurable spaces $(E,{\mathscr{E}})$ and $(M,{\mathscr{M}})$ related to the point processes given in Appendix A.
@@
For a basis $\gamma$, $t\in[0,T]$, and $\epsilon>0$, the following definitions are used in the analysis,
$\Xi^{\gamma}_{t}:=({\mathscr{W}}^{\gamma}_{s\wedge t},{\mathscr{N}}^{\gamma}_{s\wedge t})_{s\in[0,T]},\quad\Xi^{\gamma,\epsilon}_{t}:=(W^{\gamma}_{{s\vee\epsilon}\wedge t}-W^{\gamma}_{\epsilon},W^{0,\gamma}_{s\wedge t},{\mathscr{N}}^{\gamma}_{s\wedge t})_{s\in[0,T]},$ (7.1)
@@
For $\alpha^{c}\in{\mathscr{A}}(\gamma^{c}(\varrho))$, we define $\alpha^{\gamma}_{t}(\omega):=\alpha^{c}(X_{0}(\omega),{\mathscr{W}}^{\gamma}(\omega),{\mathscr{N}}^{\gamma}(\omega))$ for $t\in[0,T]$, $\omega\in\Omega^{\gamma}$. It is clear that $\alpha\in{\mathscr{A}}(\gamma)$ and ${\mathcal{L}}^{\mathbb{P}^{\gamma}}(X_{0},{\mathscr{W}}^{\gamma},{\mathscr{N}}^{\gamma},\alpha^{\gamma})={\mathcal{L}}^{\mathbb{P}^{c}}(X_{0}^{c},{\mathscr{W}}^{c},{\mathscr{N}}^{c},\alpha^{c})$. Hence, by their definitions $J^{\gamma}(X_{0},\alpha^{\gamma})=J^{\gamma^{c}(\varrho)}(X_{0}^{c},\alpha^{c})$. Since for each control in the canonical setting we have constructed a control in $\gamma$ with same cost, this implies the claimed inequality. ∎
In order to prove Theorem 7.2, in the next subsections we show that for any $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, $\gamma\in\Gamma(\rho)$, and $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$ the reverse inequality holds,
@@ 7.1 Reduction to Bounded Cost Functions
Suppose that Assumptions 2.3 and 7.1 holds. Then, for every $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, and any $c,\ell$,
@@
where we set $X^{m}:=X^{X_{0},\alpha^{m}}$, ${\mathcal{L}}^{m}:={\mathcal{L}}^{\alpha^{m}}$. By the boundedness of $\alpha^{m}$, the ${\mathscr{L}}^{2}$ boundedness of $(X^{m})_{m\geq 1}$, and basic properties of conditional expectations to deal with the ${\mathcal{L}}^{m}$-terms, it suffices to show uniform integrability of $(\|X^{m}\|_{T}^{2})_{m\geq 1}$,
@@
Using the assumptions on the SDE coefficients in Assumption 7.1 and [42, Theorem 2.11], for any fixed $\beta\in{\mathscr{A}}_{\ell}$ and $\delta>0$, there exists a constant $c$ such that
$\mathbb{E}[\|X^{m}-X^{X_{0},\beta}\|_{T}^{2+\delta}]\leq c(1+\mathbb{E}\int_{0}^{T}\|X^{m}-X^{X_{0},\beta}\|_{t}^{2+\delta}\,\mathrm{d}t),$
and we conclude by Grönwall’s inequality that $\sup_{m\geq 1}\mathbb{E}[\|X^{m}-X^{X_{0},\beta}\|_{T}^{2+\delta}]\leq ce^{cT}$. Set $A(m,K):=\{\|X^{m}\|_{T}^{2}>K\}$. Clearly, by ${\mathscr{L}}^{2}$-boundedness of $(X^{m})_{m\geq 1}$,
$\sup_{m\geq 1}\mathbb{P}(A(m,K))\leq\frac{\sup_{m\geq 1}\mathbb{E}[\|X^{m}\|_{T}^{2}]}{K}\longrightarrow 0\quad\text{as}\ K\to\infty.$
@@
$\mathbb{E}[\|X^{m}\|_{T}^{2}\chi_{A(m,K)}]$ $\leq 2\,\mathbb{E}([\|X^{m}-X^{X_{0},\beta}\|_{T}^{2}+\|X^{X_{0},\beta}\|_{T}^{2})\chi_{A(m,K)}]$
$\leq 2\,\mathbb{E}[\|X^{m}-X^{X_{0},\beta}\|_{T}^{2+\delta}]^{1/p}\mathbb{P}(A(m,K))^{1/q}+2\,\mathbb{E}[\|X^{X_{0},\beta}\|_{T}^{2}\chi_{A(m,K)}]$
$\leq 2\,(ce^{cT})^{1/p}\mathbb{P}(A(m,K))^{1/q}+2\,\mathbb{E}[\|X^{X_{0},\beta}\|_{T}^{2}\chi_{A(m,K)}].$
@@
In view of the above lemma and the discussion preceding it, this implies that $v(X_{0},\gamma)=v(X_{0}^{\prime},\gamma^{\prime})$, proving the invariance principle for the original cost functions $(F,G)$. Hence, without loss of generality we may assume that $F$, $G$ are bounded and by Assumption 7.1, there are uniformly continuous with a bounded continuous modulus of continuity $m$.

### v2

@@ 6.2 Connection to Gradient Flows
In the context of mean field control or games, an alternative connection to an optimization problem is proposed for several models, including [16, 63, 64] for the Kuramoto synchronization and [57, 58] for the Cucker-Smale flocking. In general, we propose the following MFC with a separable structure,
@@
This is again an equation of type (6.1), except the functional $F$ is replaced by the value function $v$. Now, a standard control argument shows that linear convexity of $F$ implies linear convexity of the value function $v$, suggesting a similar structure of (6.1) and (6.3). Further, note that if $F$ is sufficiently smooth, one can always define a suitable control problem such that the optimally controlled state dynamics exactly match (6.1). Indeed, this can be achieved by replacing $F$ by
$\widetilde{F}(\mu):=\beta F(\mu)+\frac{1}{2}\mu\left(|\nabla_{x}\delta_{\mu}F(\mu)(\cdot)|^{2}\right)-\frac{1}{2}\mu\left(\Delta_{x}\delta_{\mu}F(\mu)(\cdot)\right).$
Then, a direct calculation shows that the value function with running cost $\widetilde{F}$ is exactly $F$, and the optimal state process solves the mean field Langevin equation (6.1).
@@
and state dynamics $X^{\alpha}_{t}=X_{0}+\int_{0}^{t}\alpha_{s}\,\mathrm{d}s+\sigma W_{t}$. We continue to illustrating the above discussion with two prominent examples of mean field games.
@@ 6.3 Kuramoto Synchronization
An intruging property of this model is a phase transition across a critical value of $\kappa$, from incoherence to spontaneous synchronization. The related mean field model studied in [30] that shares this feature is a McKean-Vlasov equation obtained by setting $\omega^{k}$ to zero and adding a Brownian noise with strength $\sigma>0$. The resulting dynamics read,
@@
with the one-dimensional torus $\mathbb{T}$ as its state space. This flow is in the form of a mean field Langevin SDE (6.1), for the potential
@@
subject to $X^{\alpha}_{t}=\int_{0}^{t}\alpha_{s}\,\mathrm{d}s+\sigma W_{t}$. In view of Theorem 3.6, any minimizer $\alpha^{*}$ of this MFC gives rise to a mean field Nash equilibrium $\mu^{*}_{t}={\mathcal{L}}(X^{\alpha^{*}}_{t})$, in which players minimize,
@@
The above MFG approach to synchronization was first proposed by Yin, Mehta, Meyn, & Shanbhag in [64, 63], and later used by Carmona & Graves [14] to study jet-lag recovery by modeling the alignment with the circadian rhythm. These studies and the recent works of Carmona, Cormier, & Soner [16], Cesaroni & Cirant [20], and Höfer & Soner [37] establish phase transitions in the Kuramoto MFGs analogous to the one exhibited by the original and mean field Kuramoto dynamical system, providing evidence for a connection between these seemingly very different models.
@@ 6.4 Cucker-Smale Flocking
Flocking is interpreted as $\max_{k,j=1,..,N}|x^{k}_{t}-x^{j}_{t}|$ staying bounded for all times and [24] proves that flocking occurs if $\rho<1/2$. Motivated by this model, Nourian, Caines & Malhamé [57, 58] propose an ergodic mean field game in which the representative player’s running cost is given by,
@@
Consistent with this observation, [13, Chapter 4.7.3] suggests a slightly different MFG than that of [57]. It is a potential game corresponding to the MFC of (6.2) with running cost $F$ and the $2n$-dimensional state dynamics
@@ 6.5 Linear Quadratic Problem
The special case $\rho=0$ of the Cucker-Smale model is the well-known linear quadratic problem, which we now discuss. In this case, the mean field limit potential is given by,
$F(\mu)=\frac{1}{2}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}|x-y|^{2}\,\mu(\mathrm{d}x)\,\mu(\mathrm{d}y)=\frac{1}{2}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}(|x|^{2}+|y|^{2}-2x\cdot y)\,\mu(\mathrm{d}x)\mu(\mathrm{d}y)=\mathrm{Var}(\mu),$
@@
This implies that the corresponding mean field Langevin equation is given by,
@@
The stationary measures of this equation are Gaussian distributions with arbitrary mean and variance $\sigma^{2}/(4a)$. The associated potential mean field game minimizes
$J_{g,\beta}(X_{0},\alpha,\mu)=\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}|\alpha_{t}|^{2}+\kappa\int_{\mathbb{R}}|X^{\alpha}_{t}-y|^{2}\,\mu_{t}(\mathrm{d}y)\right)\,\mathrm{d}t$
for a given continuous measure flow $t\in[0,\infty)\mapsto\mu_{t}\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$. This mean field game can be solved explicitly, and a direct computation shows that any Gaussian distribution with variance $\sigma^{2}/(4a)$ provides a stationary mean field game equilibrium.
@@ 7 Invariance Principle
This section proves the invariance principle stated in Definition 3.2 in a general setting that includes law-dependent dynamics, which leads to an SDE of McKean-Vlasov type.
We fix a probabilistic basis $\gamma$ and an initial condition $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}^{\gamma}_{0})$. In this section, we allow the controlled state process to be of a general form,
$\begin{split}X^{\alpha}_{t}=X_{0}&+\int_{0}^{t}b(s,X^{\alpha}_{\cdot\land s},\alpha_{s},\mu^{\alpha}_{s})\,\mathrm{d}s+\int_{0}^{t}\boldsymbol{\sigma}(s,X^{\alpha}_{\cdot\land s},\alpha_{s},\mu^{\alpha}_{s})\,\mathrm{d}{\mathscr{W}}^{\gamma}_{s}\\ &+\int_{0}^{t}\int_{E}\boldsymbol{\lambda}(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\mu^{\alpha}_{s-},\zeta)\,\widetilde{\mathscr{N}}^{\gamma}(\mathrm{d}s,\mathrm{d}\zeta),\end{split}$ (7.1)
where $\alpha\in{\mathscr{A}}(\gamma)$, $\mu^{\alpha}_{s}:={\mathcal{L}}^{\gamma}(X^{\alpha}_{\cdot\land s}\,|\,{\mathscr{G}}^{\gamma}_{s})$, and ${\mathscr{G}}^{\gamma}$ again refers to the common noise filtration generated by $(W^{0},N^{0})$. The coefficients are given by functions,
$(b,\sigma,\sigma_{0}):[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\to\mathbb{R}^{n}\times{\mathscr{M}}_{n\times d}\times{\mathscr{M}}_{n\times\ell},$
$(\lambda,\lambda^{0}):[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\times E\to\mathbb{R}^{n}\times\mathbb{R}^{n},$
and we set $\boldsymbol{\sigma}=(\sigma,\sigma^{0})$ and $\boldsymbol{\lambda}=(\lambda,\lambda^{0})$.
The cost functional $J^{\gamma}(\alpha,X_{0})$ and value function $v(X_{0},\gamma)$ of the MFC are defined by (3.1) where $X^{\alpha}$ satisfies (7.1) instead of SDE$(X_{0},\alpha)$. Further, let $\mathfrak{m}_{2}$ be the 2-Wasserstein metric on ${\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$ and recall the sup-norm $\|\cdot\|_{T}$ on ${\mathscr{D}}_{n}$ and the notation ${\mathfrak{s}}(\nu)$ for the second moment (3.4).
The following assumption is natural extension of Assumption 2.3 and ensures strong existence and uniqueness of the general state dynamics (7.1).
@@
All coefficients are measurable and there exist constants $K,L>0$ and $K(\zeta),L(\zeta)$ such that for any $(t,x,x^{\prime},a,\mu,\mu^{\prime},\zeta)\in[0,T]\times{\mathscr{D}}_{n}^{2}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})^{2}\times E$ we have,
$|(b,\sigma,\sigma^{0})(t,x,a,\mu)-(b,\sigma,\sigma^{0})(t,x^{\prime},a,\mu^{\prime})|$ $\leq L(\|x-x^{\prime}\|_{t}+\mathfrak{m}_{2}(\mu,\mu^{\prime})),$
$|(\lambda,\lambda^{0})(t,x,a,\zeta)-(\lambda,\lambda^{0})(t,x^{\prime},a,\zeta)|$ $\leq L(\zeta)(\|x-x^{\prime}\|_{t}+\mathfrak{m}_{2}(\mu,\mu^{\prime})),$
$|(b,\sigma,\sigma^{0})(t,x,a,\mu)|^{2}$ $\leq K(1+\|x\|_{t}^{2}+|a|^{2}+\mathfrak{s}(\mu)),$
$|(\lambda,\lambda^{0})(t,x,a,\mu,\zeta)|^{2}$ $\leq K(\zeta)^{2}(1+\|x\|_{t}^{2}+|a|^{2}+\mathfrak{s}(\mu)),$
and the constants $L(\zeta),K(\zeta)$ satisfy $\int_{E}(L(\zeta)^{2}+K(\zeta)^{2})\,(n^{0}+n)(\mathrm{d}\zeta)<\infty$.
The invariance principle is proved under the following additional assumptions on the coefficients and cost functions. Set ${\mathscr{X}}:={\mathscr{D}}_{n}\times U\times{\mathscr{P}}_{2}({\mathscr{D}}_{n}\times U)$.
Assumption 7.2.
  • (Local Uniform Boundedness). For any $\ell>0$, there exists a constant $c_{\ell}$ such that for any $(t,x,\mu,\zeta)\in[0,T]\times{\mathscr{D}}_{n}\times{\mathscr{P}}_{2}({\mathscr{D}}_{n})\times E$ and $|a|,|a^{\prime}|\leq\ell$,
    $|(b,\sigma,\sigma^{0})(t,x,a,\mu)-(b,\sigma,\sigma^{0})(t,x,a^{\prime},\mu)|$ $\leq c_{\ell},$
    $|(\lambda,\lambda^{0})(t,x,a,\mu,\zeta)-(\lambda,\lambda^{0})(t,x,a^{\prime},\mu,\zeta)|$ $\leq c_{\ell}L(\zeta).$
    Further, $\int_{E}L(\zeta)^{2+\delta}(n^{0}+n)(d\zeta)<\infty$ for some $\delta>0$.
@@
  • (Uniform Continuity). For any $K>0$, there is a continuous map $m_{K}:[0,\infty)\to[0,4K]$ with $m_{K}(0)=0$ such that for all $(t,x,\hat{x},\nu,\hat{\nu},a,\hat{a})\in[0,T]\times{\mathscr{X}}^{2}$,
@@
If the coefficients of an MFC problem satisfy Assumptions 7.1 and 7.2, then it is invariant under probability bases.
@@
In the diffusion setting, the invariance is proved in [23, Theorem 3.6] for ‘rich enough’ probabilistic structures [23, Section 2.1.3] and in [29, Proposition 2.4]. The counter example provided in [23, Section 3] shows the necessity of continuity assumptions. A careful analysis of the proof of our results shows that local uniform boundedness assumption can be replaced by ${\mathscr{L}}^{p}$ integrability of $X_{0}$ for some $p>2$, a condition used in several places, for example [47], to overcome similar difficulties.
This section is devoted to the proof of Theorem 7.3. Let ${\mathscr{C}}_{m}$ be the set of $\mathbb{R}^{m}$ valued continuous functions on $[0,T]$, and recall the measurable spaces $(E,{\mathscr{E}})$ and $(M,{\mathscr{M}})$ related to the point processes given in Appendix A.
@@
For a basis $\gamma$ and $\epsilon>0$, the following definitions are used in the analysis,
$\Xi^{\gamma}_{t}:=({\mathscr{W}}^{\gamma}_{s\wedge t},{\mathscr{N}}^{\gamma}_{s\wedge t})_{s\in[0,T]},\quad\Xi^{\gamma,\epsilon}_{t}:=(W^{0,\gamma}_{s\wedge t},W^{\gamma}_{{s\vee\epsilon}\wedge t}-W^{\gamma}_{\epsilon},{\mathscr{N}}^{\gamma}_{s\wedge t})_{s\in[0,T]},\quad t\in[0,T],$ (7.2)
@@
For $\alpha^{c}\in{\mathscr{A}}(\gamma^{c}(\varrho))$, we define $\alpha^{\gamma}_{t}(\omega):=\alpha^{c}_{t}(X_{0}(\omega),{\mathscr{W}}^{\gamma}(\omega),{\mathscr{N}}^{\gamma}(\omega))$ for $t\in[0,T]$, $\omega\in\Omega^{\gamma}$. It is clear that $\alpha\in{\mathscr{A}}(\gamma)$ and ${\mathcal{L}}^{\mathbb{P}^{\gamma}}(X_{0},{\mathscr{W}}^{\gamma},{\mathscr{N}}^{\gamma},\alpha^{\gamma})={\mathcal{L}}^{\mathbb{P}^{c}}(X_{0}^{c},{\mathscr{W}}^{c},{\mathscr{N}}^{c},\alpha^{c})$. Hence, by their definitions, $J^{\gamma}(X_{0},\alpha^{\gamma})=J^{\gamma^{c}(\varrho)}(X_{0}^{c},\alpha^{c})$. Since for each control in the canonical setting we have constructed a control in $\gamma$ with same cost, this implies the claimed inequality. ∎
In order to prove Theorem 7.3, we show in the next subsections that for any $\varrho\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, $\gamma\in\Gamma(\rho)$, and $X_{0}\in{\mathcal{I}}^{\gamma}(\varrho)$ the reverse inequality holds,
@@ 7.1 Reduction to Bounded Cost Functions
Suppose that Assumptions 2.3 and 7.2 holds. Then, for every $X_{0}\in{\mathscr{L}}^{2}({\mathscr{F}}_{0})$, and any $c,\ell$,
@@
where we set $X^{m}:=X^{\alpha^{m}}$ and ${\mathcal{L}}^{m}_{t}:={\mathcal{L}}^{\alpha^{m}}_{t}={\mathcal{L}}((X^{m}_{\cdot\land t},\alpha^{m}_{t})\,|\,{\mathscr{G}})$. By the boundedness of $\alpha^{m}$, the ${\mathscr{L}}^{2}$ boundedness of $(X^{m})_{m\geq 1}$, and basic properties of conditional expectations to deal with the ${\mathcal{L}}^{m}$-terms, it suffices to show uniform integrability of $(\|X^{m}\|_{T}^{2})_{m\geq 1}$,
@@
Using the assumptions on the SDE coefficients in Assumption 7.2 and [45, Theorem 2.11], for any fixed $\beta\in{\mathscr{A}}_{\ell}$ and $\delta>0$ as in Assumption 7.2, there exists a constant $c$ such that
$\mathbb{E}[\|X^{m}-X^{\beta}\|_{T}^{2+\delta}]\leq c(1+\mathbb{E}\int_{0}^{T}\|X^{m}-X^{\beta}\|_{t}^{2+\delta}\,\mathrm{d}t),$
and we conclude by Grönwall’s inequality that $\sup_{m\geq 1}\mathbb{E}[\|X^{m}-X^{\beta}\|_{T}^{2+\delta}]\leq ce^{cT}$. Set $A(m,K):=\{\|X^{m}\|_{T}^{2}>K\}$. Clearly, by ${\mathscr{L}}^{2}$-boundedness of $(X^{m})_{m\geq 1}$,
$\sup_{m\geq 1}\mathbb{P}(A(m,K))\leq\frac{1}{K}\sup_{m\geq 1}\mathbb{E}[\|X^{m}\|_{T}^{2}]\longrightarrow 0,\quad\text{as}\ K\to\infty.$
@@
$\mathbb{E}[\|X^{m}\|_{T}^{2}\chi_{A(m,K)}]$ $\leq 2\,\mathbb{E}([\|X^{m}-X^{\beta}\|_{T}^{2}+\|X^{\beta}\|_{T}^{2})\chi_{A(m,K)}]$
$\leq 2\,\mathbb{E}[\|X^{m}-X^{\beta}\|_{T}^{2+\delta}]^{1/p}\mathbb{P}(A(m,K))^{1/q}+2\,\mathbb{E}[\|X^{\beta}\|_{T}^{2}\chi_{A(m,K)}]$
$\leq 2\,(ce^{cT})^{1/p}\mathbb{P}(A(m,K))^{1/q}+2\,\mathbb{E}[\|X^{\beta}\|_{T}^{2}\chi_{A(m,K)}].$
@@
In view of the above lemma and the discussion preceding it, this implies that $v(X_{0},\gamma)=v(X_{0}^{\prime},\gamma^{\prime})$, proving the invariance principle for the original cost functions $(F,G)$. Hence, without loss of generality we may assume that $F$, $G$ are bounded and by Assumption 7.2, they are uniformly continuous with a bounded continuous modulus of continuity $\boldsymbol{m}$.